OpenStudy (anonymous):

can somebody explain how to solve these? Solve for y. y/m+y-1/n=y-2/mn

OpenStudy (anonymous):

theyre fractions btw.

OpenStudy (loser66):

just make them have the same denominator, and then isolate y

OpenStudy (anonymous):

im confused on how to make the common denominators with letters..

OpenStudy (anonymous):

not really good with fractions period actually

OpenStudy (loser66):

\[\dfrac{y}{m}*\dfrac{n}{n}+\dfrac{(y-1)}{n}*\dfrac{m}{m}=\dfrac{(y-2)}{mn}\]